Keisha made 12 out of her last 58 shots on goal. Estimate the probability that she will make her next shot on goal.
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

12/58=.21 or there is a 21% chance that she will make her next goal based on her past performance..

Answer by stanbon(75887)

Keisha made 12 out of her last 58 shots on goal. Estimate the probability that she will make her next shot on goal.
----------------
Prob(makes next shot) = 12/58 = 6/29
=================
Cheers,
Stan H.